# Mitigating the Impact of Malware Evolution on API Sequence-based Windows Malware Detectors

### v1

@@ VII-C2 Results
Intuitively, both the embedding and encoder enhancement demonstrate significant improvements to the model, indicating that optimizing the API sequence embedding and refining the training process of the encoder through contrastive learning can effectively mitigate the impact of malware evoluation. Further observations reveal some differences in the effects of the embedding and encoder enhancements. Over the four years of testing from 2018 to 2021, for the LSTM model, the embedding enhanced LSTM shows an average decrease in FNR of 5.7% and an average increase in F1 score of 3.9% relative to the regular LSTM, while the encoder enhanced LSTM exhibited an decrease in FNR of 3.5% and an increase in F1 score of 2.1%. For the Text-CNN model, the embedding enhanced Text-CNN displays an average decrease in FNR of 8.4% compared to the regular Text-CNN, with an average increase in F1 score of 5%. Meanwhile, the encoder enhanced Text-CNN shows an average decrease in FNR of 6.4% and an average increase in F1 score of 3.3%. From these results, it appears that embedding enhancement has a slightly better effect than encoder enhancement. This suggests that a well-designed feature representation is crucial for mitigating model aging. Finally, the MME framework proposed in this paper combines both embedding and encoder enhancements and achieves the best mitigation effects. Over the four-year testing period, the MME enhanced LSTM shows an average decrease in FNR of 12.2% and an average increase in F1 score of 7.2% compared to the regular LSTM. Moreover, the MME-enhanced Text-CNN achieves an average decrease in FNR of 13.1% compared to the regular Text-CNN, with an average increase in F1 score of 8.5%.
@@ VII-D1 Experimental Settings
Here is our evaluation methodology, which involves four steps. First, we select the top 10 malware families with the most number of samples from the dataset in §VI-A. As a result, we have 17,288 malware samples in this experiment and every family has more than 1k samples. Second, for each malware family, we sort all the family samples by their appearing time and then divide them into 10 groups so that each group contains 10% samples of the family. The samples in one group is strictly ahead of samples from the next group in terms of their appearing time. Third, for each malware sample, we input its raw API sequence into the regular/MME-enhanced model and take the output of the regular/constrastive encoder as its feature representation. Lastly, we calculate a feature stability score of every two adjacent groups using Jensen–Shannon divergence [44]. The Jensen-Shannon divergence is a method used to measure the similarity between two feature distributions: . It calculates the average Kullback-Leibler divergence between the two distributions (i.e., ) and derives a final measurement value by utilizing the symmetry of the logarithmic function. In this experiment, and refer to the set of softmax-normalized features obtained in the third step for two adjacent groups. The score of ranges from 0 to 1, where the value closer to 0 indicates that the malware feature distributions between two groups are more similar, implying better feature space stability.
@@ VII-D2 Result
Figure 8 and 9 show the distribution of feature stability scores (i.e., scores) for each malware family with the regular and MME-enhanced models. We can observe that for each malware family, the JS scores of all MME-enhanced models are closer to 0, significantly lower than the results of the regular model. This indicates that the feature stability of the MME-enhanced models demonstrates better performance. During the evolution of malware, the MME enhanced model can reduce the feature space disturbances. This experiment explains why MME can help the model mitigate the impact of malware evolution, as malware tends to retain semantic similarities during its evolution, and MME can capture these similarities and maintain the feature space stability.
@@ VIII-A API Sequence-based Malware Detection
Inspired by deep learning-based squence analysis, many researchers apply some DL models like convolutional neural networks (CNNs) and recurrent neural networks (RNNs) to learn features of the API call sequences. Kolosnjaji et al. [3] use the API sequence as input. Their approach stacks a CNN that uses a 3-sized filter to represent 3 consecutive APIs (like the 3-gram approach). After the CNN, the LSTM is applied to handle the time-series sequence. Agrawal et al. [4] input the API names and the n-gram of the string arguments into several stacked LSTMs. Zhang et al. [10] build a feature engineering about the API names and arguments and then design a deep learning model including gate-CNNs and Bi-LSTM as the malware detector. Catak et al. [7] input API sequences into LSTMs to detect and classify malware. Li et al. [5] combine the Text-CNN with Bi-LSTM to analyze the API sequences and detect malware. Similarly, Chen et al. [6] use the Text-CNN and Bi-LSTM as the baseline models for API sequence-based malware detection.
Obviously, too many works have already using Text-CNN and LSTM models or their variants or combinations as the encoder of the detection models. However, how to perform feature representation so that the model can accurately understand the semantic information in API sequences and thus understand the software behavior remains a challenging issue. Our framework MME enhance the API sequence embedding and can better represent each API as a semantic feature vector, as shown in §III and §IV.
@@ IX-A Other types of Detectors
The API sequence is a popular type of feature widely adopted by dynamic malware detectors [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], mainly because API sequences are essential in understanding malware behaviors. In our experiments, we validate the effectiveness of MME by enhancing the LSTM and Text-CNN models. There indeed are many other DNN models such as Bi-LSTM, Gate-CNN, or even more variant models commonly used for learning API sequences [4, 5, 6, 10, 47, 48]. Although these models are not individually validated in this paper, they also require the API sequence embedding and encoder training process. We believe that the MME approach can also be applied to these models and achieve similar enhancement effects.
@@ IX-B Overly Advanced Malware
In reality, there are a few instances where the attack methods of certain malware are so advanced that their behavior bears very little similarity to previous malware. For such newly emerged malware, the performance of MME may weaken. In such cases, drift sample detection methods [16, 17] can be used to identify such samples, as they deviate significantly from the original data distribution. Then, such overly advanced malware can be collected and labeled for updating models. We believe that utilizing the MME enhanced models in conjunction with periodic drift sample detection can better address the continuous evolution of malicious software.
@@ IX-C Malware Sandbox Evasion
To obtain the raw API sequences, sandboxes are widely used to execute malware inside virtual machines and monitor APIs with API hooking techniques. Sandbox evasion refers to techniques employed by malware to avoid detection or analysis within a sandbox environment [49]. For example, when malware detects that it is running within a sandbox environment, it can refrain from executing any malicious operations, or even disguise itself as a legitimate application to exhibit benign behavior. The key to combating such “environment-aware” malware is to optimize the sandbox environment to closely resemble a real system environment. In the sandbox used for experiments, we simulate some user actions (such as clicking a button, typing some texts, etc.) to trigger malware real behaviors. We further utilize the statistical model proposed by Miramirkhani et al. [50] to optimize and fine-tune the sandbox, making it even closer to a real system environment. We believe these operations can help the sandbox capture the real API sequences of malware. In future works, solutions [51, 52, 53, 54] focusing on detecting sandbox evasion can be used to further optimize this issue.
@@ X Conclusion
This paper proposes a model enhancement method MME to mitigate the impact of malware evolution on API sequence-based windows malware detectors. We observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics including equivalent API usage, similar system resources, and similar API fragments. This provides an opportunity to reduce the feature gaps caused by evolution, and slowing down model aging. Firstly, by establishing an API knowledge graph and capture semantic similarities between APIs, the influence of equivalent API substitution is reduced. Secondly, by adopting hierarchical system resource encoding based on feature hashing, the model’s attention to the similarity of system resource access before and after the evolution of malware samples is enhanced. Finally, by designing a contrastive learning strategy, the model’s attention to the similar API fragments retained before and after malware evolution is strengthened. Experimental results show that MME can greatly extend the life-time of the API sequence-based malware detectors and can significantly save the human labeling efforts required for model maintenance. MME method can be applied to most API sequence-based deep learning malware detection models and help them achieve better sustainable usage.

### v2

@@ VII-C2 Results
Intuitively, both embedding and encoder enhancements have demonstrated significant improvements to the model, suggesting that optimizing API sequence embeddings through contrastive learning and refining the encoder’s training process can effectively mitigate the impact of malware evolution. Further observations reveal differences in the effects of these enhancements. Over the four-year testing period from 2018 to 2021, for the LSTM model, the embedding-enhanced LSTM model exhibited an average decrease of 5.7% in FNR and an average increase of 3.9% in F1 score compared to the regular LSTM model. In contrast, the encoder-enhanced LSTM model showed an average decrease of 3.5% in FNR and an average increase of 2.1% in F1 score. For the Text-CNN model, the embedding-enhanced Text-CNN achieved an average decrease of 8.4% in FNR and an average increase of 5% in F1 score compared to the regular Text-CNN, while the improved Text-CNN model displayed an average decrease of 6.4% in FNR and an average increase of 3.3% in F1 score. For the Transformer model, compared to the regular Transformer, our proposed MME framework model recorded an average decrease of 0.7% in FPR and an average increase of 0.32% in F1 score. Although this improvement is not substantial, it still reflects an overall enhancement. From these results, it is evident that embedding-enhanced outperforms encoder enhancement slightly, indicating that well-designed feature representations play a critical role in mitigating model aging. Finally, the proposed MME framework, which integrates both embedding and encoder enhancements, achieved the most effective mitigation. Over the four-year testing period, compared to the regular LSTM, the MME-enhanced LSTM demonstrated an average decrease of 12.24% in FNR and an average increase of 7.2% in F1 score. Similarly, compared to the regular Text-CNN, the MME-enhanced Text-CNN showed an average decrease of 13.10% in FNR and an average increase of 8.47% in F1 score. Notably, the embedding-enhanced Transformer exhibited the most significant reduction in FNR.
@@ VII-D1 Experimental Settings
Here is our evaluation methodology, which involves four steps. First, we select the top 10 malware families with the most number of samples from the dataset in §VI-A. As a result, we have 17,288 malware samples in this experiment and every family has more than 1k samples. Second, for each malware family, we sort all the family samples by their appearing time and then divide them into 10 groups so that each group contains 10% of the family. The samples in one group is strictly ahead of the samples from the next group in terms of their appearance time. Third, for each malware sample, we input its raw API sequence into the regular/MME-enhanced model and take the output of the regular/contrastive encoder as its feature representation. Lastly, we calculate a feature stability score of every two adjacent groups using Jensen–Shannon divergence [48]. The Jensen-Shannon divergence is a method used to measure the similarity between two feature distributions: . It calculates the average Kullback-Leibler divergence between the two distributions (i.e., ) and derives a final measurement value by utilizing the symmetry of the logarithmic function. In this experiment, and refer to the set of softmax-normalized features obtained in the third step for two adjacent groups. The score of ranges from 0 to 1, where the value closer to 0 indicates that the malware feature distributions between two groups are more similar, implying better feature space stability.
@@ VII-D2 Result
Figure 9, 10 and 11 show the distribution of feature stability scores (i.e., scores) for each malware family with the regular and MME-enhanced models. We can observe that for each malware family, the JS scores of all MME-enhanced models are closer to 0, significantly lower than the results of the regular model. This indicates that the feature stability of the MME-enhanced models demonstrates better performance. During the evolution of malware, the MME enhanced model can reduce the feature space disturbances. This experiment explains why MME can help the model mitigate the impact of malware evolution, as malware tends to retain semantic similarities during its evolution, and MME can capture these similarities and maintain the feature space stability.
VII-E Comparison With Other Methods
| Dataset | 2018 | 2019 | 2020 | 2021 | |||||||||||||
| Method | Acc | F1 | FPR | FNR | Acc | F1 | FPR | FNR | Acc | F1 | FPR | FNR | Acc | F1 | FPR | FNR | |
| MalScan | Degree | 76.32 | 66.67 | 3.37 | 47.98 | 73.52 | 53.44 | 3.43 | 61.64 | 44.03 | 61.14 | 0.00 | 55.97 | 49.08 | 65.84 | 0.00 | 50.92 |
| Closeness | 77.18 | 68.68 | 4.24 | 45.05 | 74.20 | 57.22 | 5.70 | 56.44 | 46.82 | 63.78 | 0.00 | 53.18 | 51.46 | 67.95 | 0.00 | 48.54 | |
| Harmonic | 76.09 | 66.98 | 4.85 | 46.73 | 72.67 | 53.24 | 5.43 | 60.72 | 44.82 | 61.90 | 0.00 | 55.18 | 51.06 | 67.60 | 0.00 | 48.94 | |
| Katz | 75.40 | 66.28 | 5.98 | 46.89 | 72.40 | 53.42 | 6.31 | 60.05 | 45.60 | 62.64 | 0.00 | 54.40 | 51.06 | 67.60 | 0.00 | 48.94 | |
| Concatenate | 76.60 | 67.99 | 4.99 | 45.42 | 73.81 | 56.07 | 5.46 | 57.80 | 46.78 | 63.74 | 0.00 | 53.22 | 53.46 | 69.67 | 0.00 | 46.54 | |
| MalBERT | - | 90.60 | 89.86 | 6.61 | 12.48 | 91.22 | 89.67 | 6.72 | 11.51 | 84.92 | 87.04 | 10.80 | 17.75 | 83.44 | 85.04 | 10.70 | 20.58 |
| Transformer | - | 95.88 | 95.70 | 1.67 | 6.56 | 95.11 | 94.23 | 3.13 | 7.23 | 91.22 | 92.53 | 4.01 | 11.75 | 89.56 | 90.56 | 3.06 | 15.51 |
| API2Vec++ | BiLSTM Attention | 81.70 | 82.04 | 20.20 | 16.40 | 87.45 | 87.41 | 11.90 | 13.20 | 81.70 | 82.45 | 22.60 | 14.00 | 83.90 | 84.32 | 18.80 | 13.40 |
| DMalNet | GNNs | 91.13 | 90.32 | 4.05 | 13.25 | 88.15 | 89.32 | 6.48 | 15.92 | 88.83 | 87.39 | 6.23 | 19.08 | 90.98 | 90.01 | 4.93 | 14.96 |
| Ours | MME(TextCNN) | 95.34 | 95.23 | 2.39 | 6.72 | 94.22 | 93.65 | 4.38 | 6.84 | 92.51 | 92.62 | 5.39 | 10.86 | 91.83 | 91.08 | 4.17 | 13.99 |
| MME(LSTM) | 92.95 | 92.65 | 5.96 | 8.04 | 92.67 | 91.61 | 7.27 | 7.37 | 89.52 | 91.10 | 10.12 | 11.06 | 89.14 | 89.55 | 8.62 | 14.12 | |
| MME(DMalNet) | 93.39 | 92.71 | 4.18 | 8.81 | 92.16 | 93.63 | 5.47 | 9.64 | 92.69 | 92.63 | 5.26 | 10.6 | 93.89 | 93.60 | 4.37 | 8.64 | |
| MME(Transformer) | 96.00 | 95.57 | 1.82 | 6.66 | 95.53 | 94.72 | 2.63 | 6.91 | 91.67 | 92.90 | 3.36 | 11.44 | 90.11 | 91.08 | 2.73 | 14.81 | |
| MME(BERT) | 95.10 | 94.74 | 2.68 | 7.34 | 94.59 | 93.65 | 3.92 | 7.38 | 90.99 | 92.36 | 4.81 | 11.63 | 89.97 | 91.02 | 3.91 | 14.24 | |
DMalNet[49] exports an API call graph from a sequence of API calls, transforming the relationships between API calls into structural information of the graph. Finally, a graph neural network was used to realize malware detection and type classification. MalBERT[50] proposes a static analysis of malicious code based on BERT[29], which uses the preprocessed features to statically analyze the source code of malware, and divides the existing malicious code into different representative malicious code categories. API2Vec++ [51] designs a Temporal Process Graph (TPG) to model inter-process behavior and a Temporal API Property Graph (TAPG) to model intra-process behavior, and develops a heuristic random walk algorithm to generate numerous paths that can capture fine-grained malicious family behaviors. MalScan[52] targets Android malware, while MME focuses on Windows malware. However, we believe that a comparison in terms of methodology is valuable, so we also take MalScan[52] as a baseline.
The experimental results, as shown in Table IX, demonstrate that our proposed MME model consistently and significantly outperforms other models across all four datasets from 2018 to 2021. Specifically, the MME(Transformer) model achieved the highest accuracy and F1 scores in 2018 and 2019, with an accuracy of 96.00% and an F1 score of 95.57% in 2018, and an accuracy of 95.53% and an F1 score of 94.72% in 2019. Furthermore, our MME(DMalNet) model exhibited outstanding performance on more recent datasets, achieving accuracies of 92.69% in 2020 and 93.89% in 2021. The effect of MME enhancing DMalNet verifies the ability of the MME framework to enhance GNN-based methods, and reflects the high adaptability and reliability of the MME framework. Compared to traditional feature extraction methods such as MalScan [52] and existing models like MalBERT [50] and DMalNet [49], our approach demonstrates clear improvements in both accuracy and F1 score. These results indicate that the proposed MME framework for malware analysis effectively enhances classification performance, reduces misclassification errors, and offers superior practical value.
Further analysis of misclassification rates among different models reveals that our proposed MME model significantly reduces both the FNR and FPR. For instance, on the 2018 dataset, our MME (Transformer) achieved a false positive rate of only 1.82% and a false negative rate of 6.66%, substantially lower than those of MalBERT [50] (FPR of 6.61%, FNR of 12.48%) and DMalNet [49] (FPR of 4.05%, FNR of 13.25%). This advantage persisted consistently over subsequent years’ datasets, demonstrating the superior stability and generalization capability of the MME model. This experiment comprehensively validates the advantages and effectiveness of our proposed MME architecture in terms of accuracy, reliability, and robustness, positioning it as a valuable method for current malware detection tasks.
VII-F Knowledge Graph Analysis
We employed T-SNE for dimensionality reduction and K-means for clustering to visualize the API entity embeddings. The reduced data was subsequently grouped to identify the primary functional categories of APIs, resulting in 20 distinct categories. From each cluster, three API entities proximate to the centroid were randomly selected as representatives. The outcomes, illustrated in Figure 12, highlight methods such as IFaxStatus::get_Receive, IFaxStatus::get_Send, and IFaxDoc::put_DisplayName, all of which pertain to fax-related operations. Observations indicate that the knowledge graph forms compact clusters, thereby facilitating the differentiation between benign and malicious software. This suggests that the knowledge graph effectively captures salient features of malware, establishing discernible classification boundaries within the feature space.
We conducted a centrality analysis on the knowledge graph and listed the different centrality measures along with their corresponding values, as presented in Table X. We identified the top two key API entities for each centrality type, leading to the following conclusions: ldap_parse_page_control and get_RealTimeBuffersLost, as well as SHCreateShellPalette, serve as both central nodes and key bridges within the network. These entities function as critical hubs in the knowledge graph, characterized by numerous direct connections and high influence, which underscores their core status within the system. The knowledge graph elucidates their semantic and structural relationships within the network, thereby aiding in the differentiation between benign and malicious software.
| API Name | |||
| Degree |
| Closeness |
| Betweenness |
| Eigenvector |
@@ VIII-A API Sequence-based Malware Detection
Inspired by deep learning-based sequence analysis, many researchers apply some DL models like convolutional neural networks (CNNs) and recurrent neural networks (RNNs) to learn features of the API call sequences. Kolosnjaji et al. [3] use the API sequence as input. Their approach stacks a CNN that uses a 3-sized filter to represent 3 consecutive APIs (like the 3-gram approach). After the CNN, the LSTM is applied to handle the time-series sequence. Agrawal et al. [4] input the API names and the n-gram of the string arguments into several stacked LSTMs. Zhang et al. [10][55],[56] build a feature engineering about the API names and arguments and then design a deep learning model including gate-CNNs and Bi-LSTM as the malware detector. Catak et al. [7] input API sequences into LSTMs to detect and classify malware. Li et al. [5] combine the Text-CNN with Bi-LSTM to analyze the API sequences and detect malware. Similarly, Chen et al. [6] use the Text-CNN and Bi-LSTM as the baseline models for API sequence-based malware detection.
Obviously, too many works have already using Text-CNN, LSTM and Transformer models or their variants or combinations as the encoder of the detection models. However, how to perform feature representation so that the model can accurately understand the semantic information in API sequences and thus understand the software behavior remains a challenging issue. Our framework MME enhance the API sequence embedding and can better represent each API as a semantic feature vector, as shown in §III and §IV.
@@ IX-A Other types of Detectors
The API sequence is a popular type of feature widely adopted by dynamic malware detectors [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], mainly because API sequences are essential in understanding malware behaviors. In our experiments, we validate the effectiveness of MME by enhancing the LSTM, Text-CNN, and other models. There indeed are many other DNN models, such as Bi-LSTM, Gate-CNN, or even more variant models commonly used for learning API sequences [4, 5, 6, 10]. Although these models are not individually validated in this paper, they also require the API sequence embedding and encoder training process. We believe that the MME approach can also be applied to these models and achieve similar enhancement effects. Current coarse-grained methods for contrastive learning have certain limitations, and future work can propose work that explores hierarchical constraints and adversarial enhancement robustness.
@@ IX-B Overly Advanced Malware
In reality, certain malware employs highly advanced attack methods, rendering their behavior nearly indistinguishable from previously known malware. For such emerging malware, the performance of MME may decline. In this scenario, drift sample detection methods[16, 17] can be utilized to identify these samples, as they significantly deviate from the original data distribution. Subsequently, these highly advanced malware samples can be collected and labeled to update the model. It is believed that integrating the MME enhancement model with periodic drift sample detection can more effectively address the ongoing evolution of malware. Furthermore, advanced evasion techniques, such as polymorphism, virtualization detection, and adversarial triggers, fall outside the experimental scope of this study. Regarding the knowledge graph, we recognize that extracting only the first verb from API description sentences may overlook other actions, potentially impacting functional similarity judgments. Future work could incorporate large language model techniques or more sophisticated NLP methods to refine this aspect. While MME’s knowledge graph and contrastive learning may offer some robustness against these threats, their effectiveness remains untested. Future research will explore MME’s performance in these scenarios by integrating adversarial training and drift detection mechanisms.
@@ IX-C Malware Sandbox Evasion
To obtain the raw API sequences, sandboxes are widely used to execute malware inside virtual machines and monitor APIs with API hooking techniques. Sandbox evasion refers to techniques employed by malware to avoid detection or analysis within a sandbox environment [57]. For example, when malware detects that it is running within a sandbox environment, it can refrain from executing any malicious operations, or even disguise itself as a legitimate application to exhibit benign behavior. The key to combating such “environment-aware” malware is to optimize the sandbox environment to closely resemble a real system environment. In the sandbox used for experiments, we simulate some user actions (such as clicking a button, typing some text, etc.) to trigger malware’s real behaviors. We further utilize the statistical model proposed by Miramirkhani et al. [58] to optimize and fine-tune the sandbox, making it even closer to a real system environment. We believe these operations can help the sandbox capture the real API sequences of malware. In future works, solutions [59, 60, 61, 62] focusing on detecting sandbox evasion can be used to further optimize this issue.
@@ X Conclusion
This paper proposes a model enhancement method, MME to mitigate the impact of malware evolution on API sequence-based windows malware detectors. We observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics, including equivalent API usage, similar system resources, and similar API fragments. This provides an opportunity to reduce the feature gaps caused by evolution, and slowing down model aging. Firstly, by establishing an API knowledge graph and capture semantic similarities between APIs, the influence of equivalent API substitution is reduced. Secondly, by adopting a hierarchical system resource encoding based on feature hashing, the model’s attention to the similarity of system resource access before and after the evolution of malware samples is enhanced. Finally, by designing a contrastive learning strategy, the model’s attention to the similar API fragments retained before and after malware evolution is strengthened. Experimental results show that MME can greatly extend the life-time of the API sequence-based malware detectors and can significantly save the human labeling efforts required for model maintenance. The MME method can be applied to most API sequence-based deep learning malware detection models and help them achieve better sustainable usage.